# Logarithmic morphisms, tangential basepoints, and little disks

### v1

@@ 2.6. Logification
[…]
@@ 2.7. Pullback
${X}$${Y}$${(\underline{X},\underline{\phi}^{*}\mathscr{M}_{Y})}$$\scriptstyle{\phi}$
[…]
@@ 2.9. Products
[…]
@@
${Z}$${X\times_{S}Y}$${X}$${Y}$${S}$$\scriptstyle{\xi}$$\scriptstyle{\phi}$$\scriptstyle{\psi}$$\scriptstyle{p_{X}}$$\scriptstyle{p_{Y}}$$\scriptstyle{a_{X}}$$\scriptstyle{a_{Y}}$
@@
Recall from Section 2.4 that virtual morphisms $*\to*_{\mathrm{log}}$ over $\mathbb{K}$ are in bijection with invertible scalars $\mathbb{K}^{\times}$. Let $X,Y=*$, equipped with the virtual morphisms to $S=*_{\mathrm{log}}$ given by scalars $\lambda,\mu\in\mathbb{K}^{\times}$, with $\lambda\neq\mu$. Given a commutative diagram of virtual morphisms log schemes
${Z}${*}{*}${*_{\mathrm{log}}}$
@@ 2.10. Continuity principle
[…]
Let $\mathbb{K}$ be a field and let $X:=\mathsf{Spec}(t^{\mathbb{N}}\to k[\varepsilon]/(\varepsilon^{2});t\mapsto\varepsilon)$, be the log scheme over $\mathbb{K}$ obtained by pulling back the standard log structure on $\mathbb{A}^{1}_{\mathrm{log}}$ to the first-order neighbourhood of 00. Note that $\varepsilon\notin\mathcal{O}_{X}^{\times}$. Hence, for any $a,b\in\mathbb{K}^{\times}$, there is a unique virtual morphism $\phi_{a,b}\colon X\to X$ such that
@@ 3.1. Tangential basepoints as virtual morphisms
[…]
Let $\underline{X}=\mathbb{A}^{1}$ with coordinate $z$, and $\underline{D}=\{0\}$. A tangential basepoint of $(\mathbb{A}^{1},\{0\})$ is either a $\mathbb{K}$-point of $\mathbb{A}^{1}\setminus\{0\}$ or a non-zero tangent vector at 00, denoted by $c\,\partial_{z}|_{0}$ with $c\in\mathbb{K}^{\times}$. ∎
@@
Virtual morphisms $*\to X$ of log schemes over $\mathbb{K}$ are in canonical bijection with tangential basepoints of $(\underline{X},\underline{D})$.
@@
According to Section 2.7, a virtual morphism $*\to X$ is the same thing as the data of a rational point $x\in\underline{X}(\mathbb{K})$, viewed as a morphism of $\mathbb{K}$-schemes $x\colon\mathsf{Spec}(\mathbb{K})\to\underline{X}$, along with a virtual morphism from * to the pullback of the log structure,
@@
which acts as the identity on $\mathbb{K}^{\times}$. This is given by a collection of $j$ non-zero scalars $(v_{1},\ldots,v_{j})\in(\mathbb{K}^{\times})^{j}$, with $v_{i}$ the image of $z_{i}$. Associating to this tuple the tangential basepoint $v_{1}\partial_{z_{1}}|_{x}+\cdots+v_{j}\partial_{z_{j}}|_{x}$ gives a bijection between virtual morphisms $*\to X$ and tangential basepoints, as claimed.
@@ 3.2.1. Examples of tangential basepoints over $\mathbb{Z}$
Let $X$ be the divisorial log scheme over $\mathbb{Z}$ associated to the pair $(\underline{X},\underline{D}):=(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$, which is the setting in which tangential basepoints were first defined [Del89]. Note that $\underline{X}\setminus\underline{D}=\mathbb{P}^{1}\setminus\{0,1,\infty\}$ does not have any $\mathbb{Z}$-points. Let $z$ be the standard coordinate on $\mathbb{A}^{1}=\mathbb{P}^{1}\setminus\{\infty\}$, and let $w=z^{-1}$ be the standard coordinate at infinity. Since ×={±$\mathbb{Z}^{\times}=\{\pm 1\}$, one sees by adapting the proof of Theorem 3.1 that $X$ has exactly two virtual $\mathbb{Z}$-points at each of the three points $0,1,\infty\in\mathbb{P}^{1}(\mathbb{Z})$, namely the “unit” tangential basepoints
±$\pm\partial_{z}|_{0}\in T_{0}\mathbb{P}^{1}_{\mathbb{Z}}\qquad\pm\partial_{z}|_{1}=\mp\partial_{w}|_{1}\in T_{1}\mathbb{P}^{1}_{\mathbb{Z}}\qquad\pm\partial_{w}|_{\infty}\in T_{\infty}\mathbb{P}^{1}_{\mathbb{Z}}.$
These form a torsor for the automorphism group $\mathrm{Aut}(X)\cong\mathfrak{S}_{3}$.
@@ 3.2.2. Examples of tangential basepoints for a non-strict normal crossing divisor
Let $X$ be the log scheme over $\mathbb{R}$ associated to the pair $(\underline{X},\underline{D})$ where $\underline{X}=\mathbb{A}^{2}_{\mathbb{R}}$ with coordinates $(u,v)$ and $\underline{D}=\{u^{2}+v^{2}=0\}$ in $\underline{X}$. It is a normal crossing divisor that only becomes strict after base change to $\mathbb{C}$ since $\underline{D}_{\mathbb{C}}=\{(u+v\mathrm{i})(u-v\mathrm{i})=0\}$, where $\mathrm{i}$ is a fixed square root of $-1$. Let us classify virtual $\mathbb{R}$-points of $X$ whose underlying point of $\underline{X}$ is $x=(0,0)$. The log structure $\mathscr{M}_{X}|_{x}$ has sections on $\mathsf{Spec}(\mathbb{R})$ and $\mathsf{Spec}(\mathbb{C})$ given by
×$\mathbb{R}^{\times}(u^{2}+v^{2})^{\mathbb{N}}\quad\mbox{ and }\quad\mathbb{C}^{\times}(u+v\mathrm{i})^{\mathbb{N}}(u-v\mathrm{i})^{\mathbb{N}}$
respectively. It is therefore not induced by a constant pre-log structure as in the proof of Theorem 3.1. A tangential basepoint for $(\underline{X},\underline{D})$ at $x$ is equivalent to a $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$-equivariant morphism of monoids ×$\mathbb{C}^{\times}(u+v\mathrm{i})^{\mathbb{N}}(u-v\mathrm{i})^{\mathbb{N}}\to\mathbb{C}^{\times}$ acting as the identity on $\mathbb{C}^{\times}$, i.e., is a complex tangential basepoint
@@ 4.2. Positive log structures
[…]
More generally, for a Deligne–Faltings log scheme $Y=(\underline{X},\underline{D},\underline{E})$ associated to a split bundle $\underline{E}=\underline{L}_{1}\oplus\cdots\oplus\underline{L}_{n}$ over a normal crossing divisor $X=(\underline{X},\underline{D})$ over $\mathbb{C}$, the virtual points $Y(\mathbb{C})$ are given by triples $(x,v,e)$ of a point $x\in\underline{X}(\mathbb{C})$, a tangential basepoint $v$ for $(\underline{X},\underline{D})$ at $x$, and an element $e\in\underline{E}^{\circ}|_{x}$. Then $\mathrm{KN}(Y)\to\mathrm{KN}(X)$ is the $(\mathsf{S}^{1})^{n}$ bundle associated to the pullback of the (×$(\mathbb{C}^{\times})^{n}$-bundle $\underline{E}^{\circ}(\mathbb{C})\to\underline{X}(\mathbb{C})$ along the blowdown map $\mathrm{KN}(X)\to\underline{X}(\mathbb{C})$. The map $Y(\mathbb{C})\to\mathrm{KN}(Y)$ is the quotient that takes a tangential basepoint $v$ to the corresponding normal rays as in the previous example, and takes an element $e\in\underline{E}^{\circ}(\mathbb{C})$ to the element in $\underline{E}^{\circ}(\mathbb{C})/(\mathbb{R}_{>0})^{n}\cong\prod_{j}\underline{L}_{j}(\mathbb{C})/\mathbb{R}_{>0}$ giving the ray spanned by $e$ in each line of the splitting. ∎
@@ 4.4. KN-points over subrings
is the quotient that identifies two virtual $B$-points whenever they differ by rescaling by a homomorphism $\mathscr{M}^{\mathrm{gp}}\to B^{\times}\cap\mathbb{R}_{>0}$. In particular, in the extreme case $B=\mathbb{Z}$, we have ×$\mathbb{Z}^{\times}\cap\mathbb{R}_{>0}=\{\pm 1\}\cap\mathbb{R}_{>0}=\{1\}$, so there is no rescaling freedom whatsoever, and hence the map $X(\mathbb{Z})\to\mathrm{KN}_{\mathbb{Z}}(X)$ is a bijection.
[…]
@@ 5.1. Kähler differentials
The sheaf of virtual log Kähler differentials is the sheaf ${{}^{\mathbf{V}}}\Omega^{1}_{X/S}$ of $\mathcal{O}_{\underline{X}}$-modules generated by $\Omega^{1}_{\underline{X}/\underline{S}}$ and formal symbols $\mathrm{dlog}{{}^{\mathbf{V}}}(f)$ satisfying the relations (23), (25), and (26).
Note that since the differentials of invertible functions generate $\Omega^{1}_{\underline{X}/\underline{S}}$, the sheaf ${{}^{\mathbf{V}}}\Omega^{1}_{X/S}$ is generated by the elements $\mathrm{dlog}{{}^{\mathbf{V}}}(f)$ for $f\in\mathscr{M}^{\mathrm{gp}}_{X}$. Consequently, we have the following.
@@
$\phi^{*}\left(\mathrm{dlog}{{}^{\mathbf{V}}}(f)\right)=\mathrm{dlog}{{}^{\mathbf{V}}}(\phi^{*}f)$
@@
[…]
For $\mathbb{K}$ a field, consider the standard log point $*_{\mathrm{log}}$ with phantom coordinate $t$, so that $\mathscr{M}_{Y}=\mathbb{K}^{\times}t^{\mathbb{N}}$. As explained in Section 2.4, a morphism $\phi\colon X\to*_{\mathrm{log}}$ in ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ is equivalent to the datum of the element $\phi^{*}(t)\in\mathscr{M}^{\mathrm{gp}}(X)$. We claim that such a morphism is automatically differentiable. Indeed, since $\mathscr{M}_{Y}$ is generated over $\mathbb{K}^{\times}$ by $t$, it suffices to verify (27) for $f=t$. But in this case, $\alpha_{*_{\mathrm{log}}}(f)=0$ so that both the left- and right-hand sides of (27) are identically zero. The same argument shows, more generally, that if $Y$ is a log scheme whose underlying scheme is a point $\underline{Y}\cong\mathsf{Spec}(\mathbb{K})$, then every virtual morphism $X\to Y$ in ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ is differentiable. ∎
@@
[…]
Let $\mathbb{K}$ be a field. In [Ach23, Construction 3.3] Achinger constructed, for every sufficiently nice log scheme $X/\mathbb{K}$ equipped with a splitting of monoids $\varepsilon\colon\mathscr{M}_{X}/\mathcal{O}_{X}^{\times}\to\mathscr{M}_{X}$, a map on forms $\varepsilon^{\circledast}\colon\Omega^{1}_{X/\mathbb{K}}\to\Omega^{1}_{\underline{X}/\mathbb{K}}$ (and a corresponding functor on flat connections and local systems) that “is not induced by a map of log schemes $\underline{X}\to X$, though it behaves as if it was”, in that it splits the canonical map $\Omega^{1}_{\underline{X}/\mathbb{K}}\to\Omega^{1}_{X/\mathbb{K}}$. It does, however, come from a virtual morphism. Namely, under the hypotheses of op. cit., the splitting $\varepsilon$ corresponds via Section 2.4 and Section 5.1 to a differentiable virtual morphism $\phi\colon\underline{X}\to X$, giving a section of the canonical projection $X\to\underline{X}$. Then $\varepsilon^{\circledast}=\phi^{*}$ is exactly the pullback on forms induced by $\phi$, and indeed the construction in op. cit. uses the reducedness (via a special case of the continuity principle) to check the consistency axiom, exactly as in the proof of Section 5.1. ∎

### v2

@@ 2.6. Logification
$\underline{\phi}^{-1}\mathcal{O}_{Y}^{\times}$$\underline{\phi}^{-1}\alpha_{Y}^{-1}(\mathcal{O}_{Y}^{\times})$$\underline{\phi}^{-1}\mathscr{M}_{Y}$$\underline{\phi}^{-1}\mathscr{M}^{\mathrm{gp}}_{Y}$$\mathcal{O}_{X}^{\times}$$\alpha_{X}^{-1}(\mathcal{O}_{X}^{\times})$$\mathscr{M}_{X}$$\mathscr{M}^{\mathrm{gp}}_{X}$$\scriptstyle{\underline{\phi}^{*}}$$\scriptstyle{\alpha_{Y}}$$\scriptstyle{\phi_{0}^{*}}$$\scriptstyle{\phi^{*}}$$\scriptstyle{\alpha_{X}}$ (10)
Remark 2.25.
@@ 2.7. Pullback
Example 2.31 (Pullbacks to divisors).
@@ 2.9. Products
$\underline{p}^{*}\mathscr{M}_{S}$$\underline{p}_{X}^{*}\mathscr{M}_{X}$$\underline{p}_{Y}^{*}\mathscr{M}_{Y}$$\mathscr{M}_{X\times_{S}Y}$$\scriptstyle{a_{X}^{*}}$$\scriptstyle{a_{Y}^{*}}$
@@
[…]
@@
Example 2.35.
Recall from Section 2.4 that virtual morphisms $\to*_{\mathrm{log}}$ over $\mathbb{K}$ are in bijection with invertible scalars $\mathbb{K}^{\times}$. Let X,Y=X,Y=*, equipped with the virtual morphisms to $S=*_{\mathrm{log}}$ given by scalars $\lambda,\mu\in\mathbb{K}^{\times}$, with $\lambda\neq\mu$. Given a commutative diagram of virtual morphisms log schemes
@@ 2.10. Continuity principle
Example 2.37.
Let $\mathbb{K}$ be a field and let $X:=\mathsf{Spec}(t^{\mathbb{N}}\to k[\varepsilon]/(\varepsilon^{2});t\mapsto\varepsilon)$, be the log scheme over $\mathbb{K}$ obtained by pulling back the standard log structure on $\mathbb{A}^{1}_{\mathrm{log}}$ to the first-order neighbourhood of 0. Note that $\varepsilon\notin\mathcal{O}_{X}^{\times}$. Hence, for any $a,b\in\mathbb{K}^{\times}$, there is a unique virtual morphism $\phi_{a,b}\colon X\to X$ such that
@@ 3.1. Tangential basepoints as virtual morphisms
Example 3.4.
Let $\underline{X}=\mathbb{A}^{1}$ with coordinate $z$, and $\underline{D}=\{0\}$. A tangential basepoint of $(\mathbb{A}^{1},\{0\})$ is either a $\mathbb{K}$-point of $\mathbb{A}^{1}\setminus\{0\}$ or a non-zero tangent vector at 0, denoted by $c\,\partial_{z}|_{0}$ with $c\in\mathbb{K}^{\times}$. ∎
@@
Virtual morphisms $\to X$ of log schemes over $\mathbb{K}$ are in canonical bijection with tangential basepoints of $(\underline{X},\underline{D})$.
@@
According to Section 2.7, a virtual morphism $\to X$ is the same thing as the data of a rational point $x\in\underline{X}(\mathbb{K})$, viewed as a morphism of $\mathbb{K}$-schemes $x\colon\mathsf{Spec}(\mathbb{K})\to\underline{X}$, along with a virtual morphism from * to the pullback of the log structure,
@@
which acts as the identity on $\mathbb{K}^{\times}$. This is given by a collection of $j$ non-zero scalars $(v_{1},\ldots,v_{j})\in(\mathbb{K}^{\times})^{j}$, with $v_{i}$ the image of $z_{i}$. Associating to this tuple the tangential basepoint $v_{1}\partial_{z_{1}}|_{x}+\cdots+v_{j}\partial_{z_{j}}|_{x}$ gives a bijection between virtual morphisms $\to X$ and tangential basepoints, as claimed.
@@ 3.2.1. Examples of tangential basepoints over $\mathbb{Z}$
Let $X$ be the divisorial log scheme over $\mathbb{Z}$ associated to the pair $(\underline{X},\underline{D}):=(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$, which is the setting in which tangential basepoints were first defined [Del89]. Note that $\underline{X}\setminus\underline{D}=\mathbb{P}^{1}\setminus\{0,1,\infty\}$ does not have any $\mathbb{Z}$-points. Let $z$ be the standard coordinate on $\mathbb{A}^{1}=\mathbb{P}^{1}\setminus\{\infty\}$, and let $w=z^{-1}$ be the standard coordinate at infinity. Since $\mathbb{Z}^{\times}=\{\pm 1\}$, one sees by adapting the proof of Theorem 3.5 that $X$ has exactly two virtual $\mathbb{Z}$-points at each of the three points $0,1,\infty\in\mathbb{P}^{1}(\mathbb{Z})$, namely the “unit” tangential basepoints
$\pm\partial_{z}|_{0}\in T_{0}\mathbb{P}^{1}_{\mathbb{Z}}\qquad\pm\partial_{z}|_{1}=\mp\partial_{w}|_{1}\in T_{1}\mathbb{P}^{1}_{\mathbb{Z}}\qquad\pm\partial_{w}|_{\infty}\in T_{\infty}\mathbb{P}^{1}_{\mathbb{Z}}.$ (15)
These form a torsor for the automorphism group $\mathrm{Aut}(X)\cong\mathfrak{S}_{3}$. The following lemma says that there are no other tangential basepoints for $X$.
Lemma 3.8.
The $6$ tangential basepoints (15) are the only virtual $\mathbb{Z}$-points of the divisorial log scheme $(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$.
Proof.
Assume that there exists a virtual morphism $s:\mathsf{Spec}(\mathbb{Z})\to(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$ whose underlying morphism $\underline{s}:\mathsf{Spec}(\mathbb{Z})\to\mathbb{P}^{1}_{\mathbb{Z}}$ is different from 0, 11, $\infty$, and denote this point by $[a:b]$ with $a,b\in\mathbb{Z}$ coprime and $a,b,a-b\neq 0$. Let $S$ be the finite set of prime numbers $p$ such that $p$ divides one of a,b,aba,b,a-b, those three cases being mutually exclusive. One easily sees that $S$ is non-empty (this is the same thing as saying that $\mathbb{P}^{1}_{\mathbb{Z}}\setminus\{0,1,\infty\}$ does not have any $\mathbb{Z}$-points). Geometrically, $S$ is the set of closed points of $\mathsf{Spec}(\mathbb{Z})$ over which $\underline{s}$ intersects one of the three sections $0,1,\infty$ of the structure morphism $\mathbb{P}^{1}_{\mathbb{Z}}\to\mathsf{Spec}(\mathbb{Z})$. Via $\underline{s}$, the divisorial log structure $(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$ pulls back to the divisorial log structure $(\mathsf{Spec}(\mathbb{Z}),S)$, and $s$ factors through a virtual morphism of log schemes $f:\mathsf{Spec}(\mathbb{Z})\to(\mathsf{Spec}(\mathbb{Z}),S)$ whose underlying morphism of schemes is the identity. Consider a prime $p\in S$, viewed as a global section of the structure sheaf of monoids for $(\mathsf{Spec}(\mathbb{Z}),S)$. The element $f^{*}(p)$ is necessarily $\pm 1$, but because $p$ is invertible outside of the closed point $p$, we necessarily have $f^{*}(p)=p$ in the localization $\mathbb{Z}_{(p)}$. This is a contradiction, and the claim follows. ∎
@@ 3.2.2. Examples of tangential basepoints for a non-strict normal crossing divisor
Let $X$ be the log scheme over $\mathbb{R}$ associated to the pair $(\underline{X},\underline{D})$ where $\underline{X}=\mathbb{A}^{2}_{\mathbb{R}}$ with coordinates $(u,v)$ and $\underline{D}=\{u^{2}+v^{2}=0\}$ in $\underline{X}$. It is a normal crossing divisor that only becomes strict after base change to $\mathbb{C}$ since $\underline{D}_{\mathbb{C}}=\{(u+v\mathrm{i})(u-v\mathrm{i})=0\}$, where $\mathrm{i}$ is a fixed square root of 1-1. Let us classify virtual $\mathbb{R}$-points of $X$ whose underlying point of $\underline{X}$ is $x=(0,0)$. The log structure $\mathscr{M}_{X}|_{x}$ has sections on $\mathsf{Spec}(\mathbb{R})$ and $\mathsf{Spec}(\mathbb{C})$ given by
[…]
respectively. It is therefore not induced by a constant pre-log structure as in the proof of Theorem 3.5. A tangential basepoint for $(\underline{X},\underline{D})$ at $x$ is equivalent to a $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$-equivariant morphism of monoids $\mathbb{C}^{\times}(u+v\mathrm{i})^{\mathbb{N}}(u-v\mathrm{i})^{\mathbb{N}}\to\mathbb{C}^{\times}$ acting as the identity on $\mathbb{C}^{\times}$, i.e., is a complex tangential basepoint
@@ 4.2. Positive log structures
Example 4.9.
More generally, for a Deligne–Faltings log scheme $Y=(\underline{X},\underline{D},\underline{E})$ associated to a split bundle $\underline{E}=\underline{L}_{1}\oplus\cdots\oplus\underline{L}_{n}$ over a normal crossing divisor $X=(\underline{X},\underline{D})$ over $\mathbb{C}$, the virtual points $Y(\mathbb{C})$ are given by triples $(x,v,e)$ of a point $x\in\underline{X}(\mathbb{C})$, a tangential basepoint $v$ for $(\underline{X},\underline{D})$ at $x$, and an element $e\in\underline{E}^{\circ}|_{x}$. Then $\mathrm{KN}(Y)\to\mathrm{KN}(X)$ is the $(\mathsf{S}^{1})^{n}$ bundle associated to the pullback of the $(\mathbb{C}^{\times})^{n}$-bundle $\underline{E}^{\circ}(\mathbb{C})\to\underline{X}(\mathbb{C})$ along the blowdown map $\mathrm{KN}(X)\to\underline{X}(\mathbb{C})$. The map $Y(\mathbb{C})\to\mathrm{KN}(Y)$ is the quotient that takes a tangential basepoint $v$ to the corresponding normal rays as in the previous example, and takes an element $e\in\underline{E}^{\circ}(\mathbb{C})$ to the element in $\underline{E}^{\circ}(\mathbb{C})/(\mathbb{R}_{>0})^{n}\cong\prod_{j}\underline{L}_{j}(\mathbb{C})/\mathbb{R}_{>0}$ giving the ray spanned by $e$ in each line of the splitting. ∎
@@ 4.4. KN-points over subrings
is the quotient that identifies two virtual $B$-points whenever they differ by rescaling by a homomorphism $\mathscr{M}^{\mathrm{gp}}\to B^{\times}\cap\mathbb{R}_{>0}$. In particular, in the extreme case $B=\mathbb{Z}$, we have $\mathbb{Z}^{\times}\cap\mathbb{R}_{>0}=\{\pm 1\}\cap\mathbb{R}_{>0}=\{1\}$, so there is no rescaling freedom whatsoever, and hence the map $X(\mathbb{Z})\to\mathrm{KN}_{\mathbb{Z}}(X)$ is a bijection.
Example 4.11.
@@ 5.1. Kähler differentials
[…]
@@
Example 5.7.
For $\mathbb{K}$ a field, consider the standard log point $_{\mathrm{log}}$ with phantom coordinate $t$, so that $\mathscr{M}_{Y}=\mathbb{K}^{\times}t^{\mathbb{N}}$. As explained in Section 2.4, a morphism $\phi\colon X\to*_{\mathrm{log}}$ in ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ is equivalent to the datum of the element $\phi^{*}(t)\in\mathscr{M}^{\mathrm{gp}}(X)$. We claim that such a morphism is automatically differentiable. Indeed, since $\mathscr{M}_{Y}$ is generated over $\mathbb{K}^{\times}$ by $t$, it suffices to verify (28) for $f=t$. But in this case, $\alpha_{*_{\mathrm{log}}}(f)=0$ so that both the left- and right-hand sides of (28) are identically zero. The same argument shows, more generally, that if $Y$ is a log scheme whose underlying scheme is a point $\underline{Y}\cong\mathsf{Spec}(\mathbb{K})$, then every virtual morphism $X\to Y$ in ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ is differentiable. ∎
@@
Example 5.10 (Splittings on forms).
Let $\mathbb{K}$ be a field. In [Ach25, Construction 3.3] Achinger constructed, for every sufficiently nice log scheme $X/\mathbb{K}$ equipped with a splitting of monoids $\varepsilon\colon\mathscr{M}_{X}/\mathcal{O}_{X}^{\times}\to\mathscr{M}_{X}$, a map on forms $\varepsilon^{\circledast}\colon\Omega^{1}_{X/\mathbb{K}}\to\Omega^{1}_{\underline{X}/\mathbb{K}}$ (and a corresponding functor on flat connections and local systems) that “is not induced by a map of log schemes $\underline{X}\to X$, though it behaves as if it was”, in that it splits the canonical map $\Omega^{1}_{\underline{X}/\mathbb{K}}\to\Omega^{1}_{X/\mathbb{K}}$. It does, however, come from a virtual morphism. Namely, under the hypotheses of op. cit., the splitting $\varepsilon$ corresponds via Section 2.4 and Section 5.1 to a differentiable virtual morphism $\phi\colon\underline{X}\to X$, giving a section of the canonical projection $X\to\underline{X}$. Then $\varepsilon^{\circledast}=\phi^{*}$ is exactly the pullback on forms induced by $\phi$, and indeed the construction in op. cit. uses the reducedness (via a special case of the continuity principle) to check the consistency axiom, exactly as in the proof of Section 5.1. ∎